# Eigenstate Thermalization Hypothesis (ETH) for off-diagonal matrix elements in integrable spin chains

### v1

@@ A.2.2 One-spin operators
From Eq. 52 we observe that the matrix elements of $E_{i}^{(11)}$, which are proportional to the matrix elements of $A$, are written as sum of determinants as in (60). Each term in the sum is the determinant of a matrix that is obtained from the Slavnov matrix (33) (or the Gaudin matrix (43) for diagonal matrix elements) by replacing a single column with a column of the matrix $P$ defined as [31]
@@
By using Eq. 60, one obtains that the off-diagonal matrix elements of one-spin operators can then be written as
$\bra{\{\mu\}}E^{(11)}\ket{\{\lambda\}}=\,\frac{\prod_{i=1}^{M}\phi\!\left(\lambda_{i}+\frac{\eta}{2}\right)}{\prod_{i=1}^{M}\phi\!\left(\mu_{i}+\frac{\eta}{2}\right)}\\ \times\frac{\det(\mathbf{S}+\mathbf{P})}{\prod_{i=1}^{M}\prod_{j<i}\phi(\lambda_{i}-\lambda_{j})\,\phi(\mu_{j}-\mu_{i})},$ (62)
@@
Having the matrix elements in the form (62), it is now possible to treat eigenstates containing strings. Indeed, in the presence of an $n$-string $\{\mu_{1},\ldots,\mu_{n}\}$, the Slavnov matrix $\mathbf{S}$ can be modified as in Eq. 40, by recursively adding the columns of the block corresponding to the string. In order to preserve the determinant, the same operations have to be performed on the matrix $\mathbf{P}$, yielding
@@
Let us now discuss the matrix elements of $E_{i}^{(21)}$. Unlike $E_{i}^{(11)}$, $E_{i}^{(21)}$ connects eigenstates with different number of particles. Now, one can check that $E_{i}^{(21)}$ is proportional to the creation operator $C(\eta/2)$. This means that the matrix elements between states |$\ket{\{\mu_{k}\}},j=1,\ldots,M$ and |$\ket{\{\lambda_{j}\}},j=1,\ldots,M+1$ is simply given by the formula
$\bra{\{\mu\}}E^{(12)}\ket{\{\lambda\}}=\frac{\Phi(\{\lambda\})}{\prod_{k=1}^{M}\phi(\mu_{k}-\eta/2)}\\ \times\frac{\det\mathbf{S^{\prime}}}{\prod\limits_{1\leq i<j\leq M+1}\phi(\lambda_{i}-\lambda_{j})\prod\limits_{1\leq l<m\leq M}\phi(\mu_{l}-\mu_{m})}$ (64)

### v2

@@ A.2.2 One-spin operators
From Eq.˜53 we observe that the matrix elements of $E_{i}^{(11)}$, which are proportional to the matrix elements of $A$, are written as sum of determinants as in (61). Each term in the sum is the determinant of a matrix that is obtained from the Slavnov matrix (34) (or the Gaudin matrix (44) for diagonal matrix elements) by replacing a single column with a column of the matrix $P$ defined as [31]
@@
By using Eq.˜61, one obtains that the off-diagonal matrix elements of one-spin operators can then be written as
@@
[…]
Having the matrix elements in the form (63), it is now possible to treat eigenstates containing strings. Indeed, in the presence of an $n$-string $\{\mu_{1},\ldots,\mu_{n}\}$, the Slavnov matrix $\mathbf{S}$ can be modified as in Eq.˜41, by recursively adding the columns of the block corresponding to the string. In order to preserve the determinant, the same operations have to be performed on the matrix $\mathbf{P}$, yielding
@@
[…]